[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO Biweekly Status Report

   Compton Observatory Science Report #161, Friday July 8, 1994
     Chris Shrader, Compton Observatory Science Support Center

       Questions or comments can be sent to the Compton SSC.
          Phone: 301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Spacecraft Status

The Observatory, scientific instruments and all spacecraft
subsystems continue to function nearly flawlessly.

                 Science Support Center Activities

Remember all those CGRO Cycle-4 Proposals you painstakingly wrote,
duplicated umpteen times, stuffed into Express-Mail envelopes and
sent off to us minutes (or less?) prior to the deadline? Well, we
haven't forgotten them either - in fact the proposals were all
evaluated and ranked on the basis of scientific merit by panels of
anonymous referees last week. These rankings, along with the
generation of the cycle-4 viewing plan which will be done next week
are the two elements which will determine which programs are
finally approved for cycle 4. Final results should be mailed out to
each proposer later this month. The observatory was heavily
oversubscribed and a broad consensus was that the quality of the
science proposed was, in most cases, quite excellent. 

                        Instrument Reports


EGRET operations were normal during this biweekly period. Delivery
of data to the GRO SSC is on schedule. Interaction with guest
investigators remains at a good level. The data reduction schedule
is slowly recovering from the time lost due to severe thunderstorms
and the resulting damage to the EGRET computers.

The current viewing period is involves the second portion of the
PSR 1951+32 observation. Analysis is well along on the deep survey
obtained earlier in phase 3; results should be presented at the AAS
meeting and will be of considerable importance to the study of AGN.
The results on the high-energy gamma rays seen from bursts,
including the spectrum, the highest energy gamma rays observed, and
the long time period of detection of the photons, continue to
reverberate through the through the community, toppling many
theories. It seems now that any theory must include the
acceleration or release of very high energy particles that interact
for over an hour. The spectrum must also be a power law and not
thermal. Combining these findings with those of BATSE severely
restricts the possibilities open to theoretical explanations.


OSSE operations are normal.

In viewing period 333 (5-12 July), the Z-axis target is Cyg X-3 (PI
team), and the X-axis targets were PSR 0611+22 (Guest Investigator
J. Cordes) and 4U 0115+63 after its recent outburst.  The Sun is
available on the scan plane, so the slewing response is enabled. 
For the last 24 hours of vp 333, we will use two detectors to
monitor the spectral state of Cyg X-1.

Data from viewing periods 212 and 213 were delivered to the Compton
GRO Science Support Center archive this week.  The targets during
period 212 were WR 140, Cyg X-1, Cyg X-3, H 1517+65.6, and Fairall
9. The targets during period 213 were the Crab, III Zw 2, and Mrk

Preprints from a number of papers by the OSSE team were distributed
last week. If you would like to receive preprints and are not
already on our distribution list, contact Tina Obrebski
(tobrebski@osse.nrl.navy.mil, or 202-767-3165).


BATSE contribution to the GROSSC biweekly report:

Over the last week the BATSE pulsed source monitor has been
detecting pulsed flux from Cen X-3, OAO 1657-415, Vela X-1 and GX

BATSE folded-on-board data is currently being collected for PSR
0344+54, PSR 0329+54, the Crab Pulsar, and, PSR J2043+27. Single
sweep data is also being collect for pulsar searches.

On June 30th an GOES class M2.5 solar flare triggered the batse
burst mode. The flare lasted approximately 370 s and was seen above
300 keV. The last M Class flare to trigger BATSE occurred on
January 29th.

A very intense burst occurred on July 3 at 4:40 UT. The peak BATSE
count rate exceeded 100000 cnt/s.  The burst lasted approximately
90 s. The preliminary burst location is ra = 143 degrees, dec = 27
degrees (J2000), with an error circle of about 3 degrees radius.

As of July 5, BATSE has detected 1037 cosmic gamma-ray bursts out
of a total of 2956 on-board triggers in 1169 days of operation.
There have been 734 triggers due to solar flares with emission
above 60 keV.


The COMPTEL instrument is performing well and continues routine

Papers by the collaboration recently accepted for publication
include reports on "Diffuse Continuum Gamma Rays from the Galaxy
Observed by COMPTEL" (Strong et al.), and "COMPTEL Gamma-ray
Observations of the Quasars CTA 102 and 3C 454.3" (Blom et al.),
both to appear in Astronomy and Astrophysics. Preprints will be
available and distributed in the coming weeks.

A first announcement of the Second Waterville Valley Workshop on
High-energy Solar Phenomena ("The Lessons of Cycle 22 and the
Promise of Cycle 23"), to be held from March 6-10, 1995 in
Waterville Valley, New Hampshire, was recently distributed
electronically.  Potential attendees and other interested parties
should refer to the full announcement kept on the GRONEWS bulletin
board (under "GRO-related meetings").

Finally, now that all interest in the World Cup has evaporated
since Team USA was eliminated by Brazil, attention within the team
is now focused on the opening stages of the 81st Tour de France,
where the field is considered wide open in this new post-Fignon